OpenStudy (greatlife44):

Can someone help with this? cos(x)sin(y^{2}) = 3x

OpenStudy (greatlife44):

Implicit differentiation need to find \[\frac{ dy }{ dx }\]

OpenStudy (greatlife44):

\[\frac{ dy }{ dx } Cos(x)Sin(y^{2}) = \frac{ dy }{ dx }3x \]

OpenStudy (greatlife44):

Doing the right side is easy it's just \[\frac{ dy }{ dx } 3x = 3 \]

OpenStudy (greatlife44):

@Michele_Laino @imqwerty

hartnn (hartnn):

product and chain rule on left side. know these?

OpenStudy (greatlife44):

familiar with chain rule not product rule